OpenStudy (anonymous):

Solve for x: x + = 15 x = −3 x = −1.5 x = 0 x = 3

OpenStudy (anonymous):

x + what?

OpenStudy (anonymous):

x + what??

OpenStudy (anonymous):

5 over 2x plus fifteen over 2

OpenStudy (hba):

\[x+\frac{ 5 }{ 2x }=15\]

OpenStudy (anonymous):

5/2x+15/2 = 15x = −3x = −1.5x = 0x = 3 like that?

OpenStudy (anonymous):

yea

OpenStudy (hba):

Well You Dont Need To Really Need To Actually Solve It.Since You Have Answer Options You Can Put The Value And Check And When LHS And RHS Match Thats Your Answer

OpenStudy (anonymous):

there is no answer....

OpenStudy (anonymous):

at least none that i can find. but im pretty sure that @hba's answer is wrong... if you solve: x+5/2x=15 x+5=30x x=30x-5 x=0.16666666666666666666666666666667 unless that's an answer choice, idk...

OpenStudy (hba):

\[x+5/2x+15/2=15\]

OpenStudy (hba):

Btw We Dont Know What The Correct Question Is ?

OpenStudy (anonymous):

@rhondaycole can you rewrite the question using the equation button

OpenStudy (anonymous):

@rhondaycole \[x + \frac{ 5 }{ 2x } = 15\]Multiply by 5.\[x + \frac{ 5 }{ 2x }(5) = 15(5)\]\[x + 2x = 75\]x + 2x = 3x SO: 3x = 75 Divide both sides by 3. x = 25

OpenStudy (anonymous):

@surdawi @lharrell97 this right?

OpenStudy (anonymous):

@jazy not sure about the equation but you did a mistake at the beginning x + 5/2x = 15 when multiply by 5 it will be 5x + 25/2x = 75

OpenStudy (anonymous):

@jazy that doesn't work.

OpenStudy (anonymous):

You would multiply the denominator to get rid of the fraction, not the numerator.

OpenStudy (anonymous):

You would want to multiply by 2x to get it out of the denominator.

OpenStudy (anonymous):

Oh yeah! I just realized that sorry! :/

OpenStudy (anonymous):

\[x + \frac{ 5 }{ 2x } =15\] \[(2x)*x + (2x)*\frac{ 5 }{ 2x } =(2x)*15\] \[2x^{2} + 5 = 30x\] \[2x^{2} -30 x +5 = 0\] can you try now @jazy

OpenStudy (anonymous):

but still we are not sure what is we got the question right